solve each equation for the variable show your work
(c-2)(c-4)=0

Answer by jim_thompson5910(35256)

(c-2)(c-4)=0

c-2=0 or c-4=0 ... Use the zero product property here

c=2 or c=4

Answer: The solutions are c=2 or c=4
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------